OpenStudy (anonymous):

2= √(x^2-6x+3) find x

OpenStudy (earthcitizen):

\[2= √(x^2-6x+3) => (x^2-6x+3)=2^2 =>(x^2-6x+3)=4\] \[(x^2-6x+3)-4=0 => (x^2-6x+3-4)=(x^2-6x-1)\]

OpenStudy (earthcitizen):

\[(x^2-6x-1) \] use quadratic formula..

OpenStudy (earthcitizen):

\[x = 3-\sqrt(10), 3+\sqrt(10)\]

OpenStudy (anonymous):

i got 3±√(20)

OpenStudy (anonymous):

NVM ! i forgot to simplify the radical